OpenStudy (anonymous):

THE ORDERED PAIR (2,25) IS A SOLUTION FOR WHICH EQUATION(S) CHECK ALL THAT APPLY A) Y=(X+2)^2 B) Y=2X^+5 C) Y=2X=1)^2 D) Y= (X=+3^2) HELP NEEDED.

OpenStudy (anonymous):

A (2 + 3)² = 5² = 25 C (2•2 + 1)² = 5² = 25

OpenStudy (anonymous):

i mean D

OpenStudy (anonymous):

D and C

OpenStudy (anonymous):

SO A & D IS THE CORRECT ANSWER?

OpenStudy (anonymous):

A. y = (x + 3)2 C. y = (2x + 1)2

OpenStudy (anonymous):

THANKSS (:

OpenStudy (anonymous):

yw